OpenStudy (anonymous):

HELP! WILL GIVE MEDALS! Consider two lines that form a linear system. Line 1: slope = 2 and y-intercept of (0,-5) Line 2: slope = 4 and y-intercept of (0,8) Part 1: What type of lines would this be? Part 2: How many solutions does this system have?

OpenStudy (anonymous):

@mathmath333 @mathgeek27 @mathdumdum123

OpenStudy (mathmath333):

first find the equations of the two lines by formula y-y1=m(x-x1) graph these two lines if they they intersect then they are intersecting lines and have one solution if they are parallel then they are parallel lines and there is no solution

OpenStudy (anonymous):

Im really confused... @mathmath333

OpenStudy (mathmath333):

*if they dont intersect then they are parallel lines and there is no solution

OpenStudy (anonymous):

do they intersect? @mathmath333

OpenStudy (mathmath333):

the equation will tell that,have u got the equation

OpenStudy (mathmath333):

for first equation put (0,-5) ,m(slope)=2 in equation y-y1=m(x-x1)

OpenStudy (anonymous):

I dont have the equation @mathmath333

OpenStudy (mathmath333):

but u can find it ,can't u ? in the previous comment i said for first equation put (0,-5) ,m(slope)=2 in equation y-y1=m(x-x1)

OpenStudy (mathmath333):

y-(-5)=2(x-0) y+5=2x 2x-y=5.....this is the first equation like these find second also

OpenStudy (anonymous):

Im so bad at math. I dont know how to find the second equation. @mathmath333